OpenStudy (anonymous):

Solve 9x^2 - 100 = 0

OpenStudy (anonymous):

I'll start you off: \(9x^2 - 100 = 0 \Rightarrow x^2 = \dfrac{100}{9}\).

OpenStudy (anonymous):

a^2-b^2= (a-b)(a+b)

OpenStudy (anonymous):

^That is a valid technique as well.

OpenStudy (anonymous):

So what would I do after 100/9?

Directrix (directrix):

@Brookebrookiee If you are going with x^2 = 100/9, then take square roots of both sides. I usually go this route: a^2-b^2= (a-b)(a+b) 9x^2 - 100 = (3x + ?)*(3x - ?) = 0 (3x + ?)*(3x - ?) = 0 After factoring, set each factor to zero and solve for x.